# Supp_Appendix_v4

### v3

@@ Appendix E Proofs for Section 5
The first equality follows from $Y=Y(0)+D(Y(1)-Y(0))$ and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
@@ Appendix G Proofs for Appendix C
$\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right\}\cap\left\{\widehat{w}_{0}(x_{j})>c_{n}\right\}\right)$
@@
[…]
@@
$\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})=0\right\}\cup\left\{\widehat{w}_{0}(x_{j})\leq c_{n}\right\}\right)$
@@
[…]
@@
[…]
@@
[…]
@@
The third equality follows from $k\in\Psi_{\mathcal{X}^{+}}$. By the proof of Theorem C.1, $\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))=o_{p}(1)$. Since $-\sqrt{n}\xi_{n}\rightarrow-\infty$, $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})\to\mathbb{P}(0\geq-\infty)=1$ when $k\in\Psi_{\mathcal{X}^{+}}$.
@@
[…]
@@ Appendix H Additional Derivations for Difference-in-Differences
[…]

### v4

@@ Appendix E Proofs for Section 5
The first equality follows from Y=Y(0)+D(Y(1)Y(0))Y=Y(0)+D(Y(1)-Y(0)) and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
@@ Appendix G Proofs for Appendix C
$\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})$ $=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right\}\cap\left\{\widehat{w}_{0}(x_{j})>c_{n}\right\}\right)$
@@
We have that $\mathbb{P}(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0)\rightarrow 1$ since $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})\xrightarrow{p}p_{j}>0$. We also have that $\mathbb{P}(\widehat{w}_{0}(x_{j})>c_{n})=\mathbb{P}(\widehat{w}_{0}(x_{j})-c_{n}>0)\rightarrow 1$ because $\widehat{w}_{0}(x_{j})-c_{n}\xrightarrow{p}w_{0}(x_{j})>0$ by $c_{n}=o(1)$ and $w_{0}(x_{j})>0$, which follows from $x_{j}\in\mathcal{X}^{+}$. Therefore, $\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})=1-\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})\leq 1-(\mathbb{P}(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0)+\mathbb{P}(\widehat{w}_{0}(x_{j})>c_{n})-1)\rightarrow 0$ as $n\rightarrow\infty$.
@@
$\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})$ $=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})=0\right\}\cup\left\{\widehat{w}_{0}(x_{j})\leq c_{n}\right\}\right)$
@@
By Assumption C.1, $\sqrt{n}\widehat{w}_{0}(x_{j})=\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))\xrightarrow{d}\mathbb{Z}_{\textbf{w}_{0}}(j)=O_{p}(1)$, since $w_{0}(x_{j})=0$ for $x_{j}\notin\mathcal{X}^{+}$. Also $\sqrt{n}c_{n}\rightarrow+\infty$ by the theorem’s assumption. Therefore, $\mathbb{P}(\sqrt{n}\widehat{w}_{0}(x_{j})\leq\sqrt{n}c_{n})\rightarrow 1$ and $\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})\rightarrow 0$ as $n\rightarrow\infty$. Because of this,
@@
By the consistency of $\widehat{\textbf{a}}$ for a, the continuity of the maximum operator, and the continuous mapping theorem, $\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\xrightarrow{p}\max_{x\in\mathcal{X}^{+}}a(x)$. Because $\mathcal{X}^{+}=\text{supp}(X\mid W_{0}=1)$ is a finite set, we also have that $\max_{x\in\mathcal{X}^{+}}a(x)=\sup(\text{supp}(a(X)\mid W_{0}=1))$. Another application of the continuous mapping theorem suffices to show that $\widehat{\overline{P}}$ is consistent for $\overline{P}$.
@@
We verify the validity of the bootstrap by appealing to Theorem 3.2 in Fang and Santos (2019). We show that their Assumption 4 holds by showing the mapping $\widehat{\psi}$ satisfies $|\widehat{\psi}(h^{\prime})-\widehat{\psi}(h)|\leq C_{n}\|h^{\prime}-h\|$ for any $h^{\prime},h\in\mathbb{R}^{3K}$ and for $C_{n}=O_{p}(1)$, and by showing that $\widehat{\psi}(h)\xrightarrow{p}\psi(h)$ for all $h\in\mathbb{R}^{3K}$.
@@
The third equality follows from $k\in\Psi_{\mathcal{X}^{+}}$. By the proof of Theorem C.1, $\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))=o_{p}(1)$. Since $\sqrt{n}\xi_{n}\rightarrow-\infty$, $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})\to\mathbb{P}(0\geq-\infty)=1$ when $k\in\Psi_{\mathcal{X}^{+}}$.
@@
where the last equality holds from $k\notin\Psi_{\mathcal{X}^{+}}$. Therefore, $\mathbb{P}(\widehat{\Psi}_{\mathcal{X}^{+}}=\Psi_{\mathcal{X}^{+}})\rightarrow 1$ as $n\rightarrow\infty$. This implies $\widehat{\psi}(h)\xrightarrow{p}\psi(h)$, which concludes the proof. ∎
@@ Appendix H Additional Derivations for Difference-in-Differences
Here $A^{\perp B}$ is used to denote the residual in the linear projection of $A$ on $(1,B)$. It is also the case that $\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\sum_{l>k}\big(\sigma_{kl}^{k}+\sigma_{kl}^{l}\big)=1$. [^9] The result in Goodman-Bacon (2021) technically also includes a weight $\sigma_{kU}$ attached to the contrast between group $k$ and the never-treated group. We subsume this weight under $\sigma_{kj}^{k}$, and likewise subsume the weight on the contrast with the always-treated group under $\sigma_{jk}^{k}$. When we compare this representation with Proposition 5.4, that is,